# : Towards Semantic Steganography via Large Language Models

### v1

@@ A.1 IVT-S
The IVT-S dataset covers a wide range of short textual types, including real person names [name_country_origin_dataset], news headlines [ao-etal-2021-pens], paper titles [farhangi2022protoformer], Twitter comments [go2009twitter], and dialogue snippets [southparkdata]. For more details on IVT-S, please refer to Table 4.
| Category | Source |
License | ||||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 9.7 | 592.2 | 4576 | 1000 | GPL-3.0 | ||||||||
| news title | PENS [ao-etal-2021-pens] | 10.3 | 506.9 | 5506 | 1000 | MSR License | ||||||||
| article title | examiner [examine_examiner_dataset] | 8.6 | 423.1 | 4873 | 1000 | CC0 1.0 | ||||||||
| people name | name&code [name_country_origin_dataset] | 2.1 | 113.8 | 857 | 1000 | - | ||||||||
| sentence | cmn_en [manythings_anki_dataset] | 6.2 | 247.3 | 1946 | 1000 | CC-BY 2.0 | ||||||||
| dialogue | South Park [southparkdata] | 4.5 | 202.9 | 1924 | 1000 | - | ||||||||
| comment | sentiment140 [go2009twitter] | 10.7 | 486.2 | 4374 | 1000 | - | ||||||||
| overall | - | 7.4 | 367.5 | 19941 | 7000 | - |
A.2 IVT-M
@@
Messages in IVT-M are sourced from a variety of domains, including CNN news abstracts [see-etal-2017-get], Wikipedia summaries [wikidump], IMDB reviews [imdb_reviews_dataset]. For more details about IVT-M, please refer to Table 5.
| Category | Source |
License | ||||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 22.0 | 1210.0 | 7776 | 1000 | GPL-3.0 | ||||||||
| new title | PENS [ao-etal-2021-pens] | 23.0 | 1369.4 | 8003 | 1000 | MSR License | ||||||||
| news highlight | CNN and Daily mail [see2017get] | 50.3 | 2303.6 | 13119 | 1000 | Apache 2.0 | ||||||||
| translate text | WMT-14 [bojar-EtAl:2014:W14-33] | 33.8 | 1610.1 | 6760 | 1000 | - | ||||||||
| joke story | stupid joke [pungas] | 47.6 | 2042.1 | 10637 | 1000 | - | ||||||||
| translate text | europarl [koehn2005europarl] | 34.0 | 1618.8 | 6735 | 1000 | - | ||||||||
| dialogue | South Park [southparkdata] | 31.7 | 1367.1 | 3896 | 1000 | - | ||||||||
| comment | hate text [hateoffensive] | 23.4 | 989.9 | 6564 | 1000 | MIT | ||||||||
| moral | moral story [guan2022corpus] | 32.5 | 1369.4 | 3179 | 1000 | CC BY-NC-SA 3.0 | ||||||||
| overall | - | 33.1 | 1542.3 | 42923 | 9000 | - |
A.3 IVT-L
@@
IVT-L represents the largest and most challenging subset of the IVT dataset, designed to evaluate a modelβs ability to embed and recover long-form semantic content. Each secret message in IVT-L consists of a full paragraph containing multiple sentences with coherent structure, complex syntax, and a diverse vocabulary. Compared to IVT-S and IVT-M, IVT-L significantly enhances the semantic richness and information density of the hidden text, thereby imposing greater demands on a modelβs capacity, fidelity, and ability to preserve semantics. This subset serves as a benchmark for evaluating whether steganographic models can scale to real-world scenarios that involve concealing full-length, meaningful textual content.
@@
IVT-L includes long-form messages drawn from a variety of sources, including narrative news passages, short stories [guan2022corpus, eldan2023tinystories], research paper abstracts [farhangi2022protoformer], IMDB reviews [imdb_reviews_dataset]. These samples span multiple domains and writing styles, providing a comprehensive benchmark for evaluating sentence-to-image steganography in high-capacity settings. For more details about IVT-L, please refer to Table 6.
| Category | Source |
License | ||||||||
| paper abstract | arxiv-10 [farhangi2022protoformer] | 156.55 | 8512.75 | 22550 | 1000 | GPL-3.0 | ||||||||
| wiki abstract | wiki [wikidump] | 140.67 | 6962.01 | 25117 | 1000 | GFDL/CC BY-SA | ||||||||
| story | stupid joke [pungas] | 153.37 | 6725.4 | 23701 | 1000 | - | ||||||||
| story | moral story [guan2022corpus] | 225.92 | 9679.95 | 18412 | 1000 | - | ||||||||
| story | Tiny Story [eldan2023tinystories] | 172.33 | 7073.81 | 10477 | 1000 | - | ||||||||
| user review | IMDB [imdb_reviews_dataset] | 194.06 | 8761.54 | 30569 | 1000 | IMDB License | ||||||||
| user review | yelp [asghar2016yelp] | 128.16 | 5542.79 | 17244 | 1000 | Yelp License | ||||||||
| essay | asap-aes [asap_aes_dataset] | 199.43 | 8734.33 | 15142 | 1000 | - | ||||||||
| overall | - | 171.3113 | 7749.073 | 111662 | 8000 | - |
Appendix B Experimental Details
@@
To explore the feasibility and performance of sentence-to-image steganography powered by large language models, we propose three variants of our framework, namely -Qwen2.5-0.5B, -Llama3.2-1B, and -MiniCPM-1B. These models are built upon different backbone LLMs with varying scales and capabilities, providing diverse perspectives on model behavior across capacity and architecture. In the following sections, we detail the training procedures and dataset construction strategies used for each model, including how the secret messages are formatted, encoded, and paired with cover images to form effective training samples. This enables a thorough evaluation of how different LLMs perform under the semantic steganography task.
B.1 Implementation Details
@@
-Qwen2.5-0.5B. -Qwen2.5-0.5B is built upon the Qwen2.5-0.5B backbone [qwen2025qwen25technicalreport] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
-Llama3.2-1B. -Llama3.2-1B is built upon the Llama3.2-1B backbone [grattafiori2024llama] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
-MiniCPM-1B. -MiniCPM-1B is built upon the MiniCPM-1B backbone [hu2024minicpm] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 24 and train for one epoch. In Stage 2, we train for 5,000 iterations with a batch size of 24.
@@
-Gemma3-1B. -Gemma3-1B is built upon the Gemma3-1B backbone [team2025gemma] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
StegaStamp. We retrain StegaStamp [tancik2020stegastamp] with extended capacity (from 100 to 512 bits) on cover images with 256Γ256 resolution using the Flickr2K dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 15,000 iterations with batch size set to 16. While achieving stable convergence, we observe fundamental limitations: the 512-bit encoding yields only ~70% bit accuracy, resulting terrible performance in IVT. This suggests current spatial-domain watermarking architectures struggle with high-capacity payloads with primitive representation of secret message such as UTF codes.
@@
LanNet. We retrain LanNet [lan2023robust] on cover images with 256Γ256 resolution using the DIV2K training dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 2,000 epochs with batch size set to 16.
B.2 Details about Training Dataset
To construct the training dataset for the aforementioned models, we curated a diverse collection of textual data from multiple sources. Specifically, we selected 14,000 highlights from the CNN/DailyMail dataset [see-etal-2017-get], 5,000 news headlines from the PENS dataset [ao-etal-2021-pens], 5,000 scientific titles from the ArXiv-10 dataset [farhangi2022protoformer], and 4,000 story paragraphs from the CNN Story corpus [cnn_stories_dataset]. This composition ensures a broad range of semantic structures and topics, facilitating robust model training across different text granularities.
Importantly, we ensure that there is no overlap between the training set and the proposed IVT benchmark datasets, thereby preventing data leakage and preserving the integrity of the evaluation. For more details about IVT-L, please refer to Table 7.
| Category | Source |
| news headlines | PENS [ao-etal-2021-pens] | 21.6 | 1296.3 | 5000 | ||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 25.1 | 1402.0 | 5000 | ||||||
| news highline | CNN/DailyMail [see-etal-2017-get] | 55.8 | 2826.4 | 14000 | ||||||
| new story | CNN Story [cnn_stories_dataset] | 362.6 | 11024.6 | 4000 |
Appendix C Steganalysis
C.1 Statistical Steganalysis
@@
We use StegExpose [Boehm2014StegExposeA] to measure the anti-steganalysis ability. The receiver operating characteristic (ROC) curve in FigureΒ 8 shows that has high security and can fool the StegExpose.
Appendix D VLMs in
In the previous experiments, all models were based on pure text-only large language models. To further explore the applicability of vision-language models (VLMs) to the sentence-to-image steganography task, we conducted additional experiments using Qwen2-vl [Qwen2-VL] as a representative VLM. However, the results showed that Qwen2-vl failed to fit the training data.
We hypothesize that this underperformance stems from the fundamental difference between visual understanding and steganography. While VLMs are primarily optimized for visual tasks, which focusing on understanding objects, scenes, and spatial relationships within images. However, steganography emphasizes decoding secret messages by focusing on subtle, imperceptible signals embedded within images. Unlike typical vision-language alignment tasks, steganography does not require high-level visual semantics but rather the capacity to detect and interpret fine-grained patterns that are deliberately encoded. This highlights the need for models specifically tailored to the semantic embedding and recovery processes intrinsic to image-based steganography.
In experiments, the -Qwen2-vl-2B model can output the and tokens, but the recovery information is totally different from the ground truth. That may because the knowledge in the vision encoder seriously hindering the language model to get exact secret information.
Appendix E Experiments on Generative Message
To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM. Specifically, we leverage the DeepSeek [deepseek2024] API to generate the textual content, as illustrated in Appendix Section F. Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
Appendix F Codes for Collecting IVTG

### v2

@@ 7.1 IVT-S
7.2 IVT-M
@@
7.3 IVT-L
@@
IVT-L represents the largest and most challenging subset of the IVT dataset, designed to evaluate a modelβs ability to embed and recover long-form semantic content. Each secret message in IVT-L consists of a full paragraph containing multiple sentences with coherent structure, complex syntax, and a diverse vocabulary, with a length of about 100 to 500 words. Compared to IVT-S and IVT-M, IVT-L significantly enhances the semantic richness and information density of the hidden text, thereby imposing greater demands on a modelβs capacity, fidelity, and ability to preserve semantics. This subset serves as a benchmark for evaluating whether steganographic models can scale to real-world scenarios that involve concealing full-length, meaningful textual content.
@@
IVT-L includes long-form messages drawn from a variety of sources, including narrative news passages, short stories [22, 15], research paper abstracts [16], IMDB reviews [49]. These samples span multiple domains and writing styles, providing a comprehensive benchmark for evaluating sentence-to-image steganography in high-capacity settings. For more details about IVT-L, please refer to Table 6.
7.4 IVTG
To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM. Specifically, we leverage the DeepSeek [14] API to generate the textual content. Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
8 Experimental Details
@@
To explore the feasibility and performance of sentence-to-image steganography powered by large language models, we propose three variants of our framework, namely -Qwen2.5-0.5B, -Llama3.2-1B, and -MiniCPM-1B. These models are built upon different backbone LLMs with varying scales and capabilities, providing diverse perspectives on model behavior across capacity and architecture. In the following sections, we detail the training procedures and dataset construction strategies used for each model, including how the secret messages are formatted, encoded, and paired with cover images to form effective training samples. This enables a thorough evaluation of how different LLMs perform under the semantic steganography task. For all baseline methods, we follow the training settings described in their original papers to ensure a fair comparison.
8.1 Training Dataset of
Text Dataset. We use the English subset of the WanJuan [25] Text Dataset as our text dataset. WanJuan 1.0 Text Dataset is composed of cleaned pre-training corpora from different sources such as web pages, encyclopedias, books, patents, textbooks, and exam questions. The total amount of data exceeds 500 million documents, and the data size exceeds 1TB. We selected a subset of the English dataset (4 GB). For texts that spanned multiple lines, we removed ββ, ββ and extra spaces to merge them into single lines. We then extracted texts with fewer than 1,000 words, resulting in 3,537,390 samples. Finally, we randomly selected 320,000 samples to construct the training set. The distribution of text lengths in the training set is shown in Fig.Β 7.
Image Dataset. We use the COCO Training Dataset [41] training dataset as our cover image dataset.
8.2 Implementation Details
@@
-Qwen2.5-0.5B. -Qwen2.5-0.5B is built upon the Qwen2.5-0.5B backbone [55] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
-Llama3.2-1B. -Llama3.2-1B is built upon the Llama3.2-1B backbone [21] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
-MiniCPM-1B. -MiniCPM-1B is built upon the MiniCPM-1B backbone [28] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 24 and train for one epoch. In Stage 2, we train for 5,000 iterations with a batch size of 24.
@@
-Gemma3-1B. -Gemma3-1B is built upon the Gemma3-1B backbone [61] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
@@
StegaStamp. We retrain StegaStamp [60] with extended capacity (from 100 bits to 512 bits) on cover images with resolution using the Flickr2K dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 15,000 iterations with batch size set to 16. While achieving stable convergence, we observe fundamental limitations: the 512-bit encoding yields only 70% bit accuracy, resulting terrible performance in IVT. This suggests current spatial-domain watermarking architectures struggle with high-capacity payloads with primitive representation of secret message such as UTF codes. More details can be found in Sec.Β 12.2.
@@
LanNet. We retrain LanNet [36] on cover images with resolution using the DIV2K training dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 1,000 epochs with batch size set to 12.
FPGP. We retrain FPGP [78] on cover images with resolution using the DIV2K training dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 1,000 epochs with batch size set to 12.
CRMark. We retrain CRMark [10] on cover images with resolution using the DIV2K training dataset. The model is optimized with AdamW with learning rate set to 1e-4 for 1000 epochs with batch size set to 12.
| Methods | -Q | DwtDct | StegaStamp | LanNet | FPGP | CRMark |
| K-L | 0.0015 | 0.082 | 0.072 | 0.0024 | 0.0017 | 0.0019 |
9 Statistical Steganalysis
@@
We use StegExpose [6] to measure the anti-steganalysis ability. The receiver operating characteristic (ROC) curve in FigureΒ 9 shows that has high security and can fool the StegExpose.
K-L divergence provides a convenient means of gauging how easy it is to discriminate between cover and stego. Typical cover data, however, are not i.i.d. (independent and identically distributed). For example, both pixels and audio samples are known to be highly correlated [58]. Therefore, KL divergence is insufficient as a standalone criterion for evaluating steganographic security [58]. Nevertheless, we compute the K-L divergence between the stego images generated by and the original cover images. As shown in Tab.Β 7, all steganographic models achieve very small KL divergence values, with getting the best result.